OpenStudy (destinyyyy):

Find the coordinates of the vertex and the intercepts of the following quadratic function. When necessary, approximate the x-intercepts to the nearest tenth.

OpenStudy (destinyyyy):

f(x)+x^2+6x+5

OpenStudy (destinyyyy):

a=1 b=6 c=5 =-(6)/2(1)

OpenStudy (princeharryyy):

y = (x+1)(x+5), I guess you can solve the rest by yourself

OpenStudy (destinyyyy):

??

OpenStudy (princeharryyy):

f(x)+x^2+6x+5 can be expressed as y = (x+1)(x+5)

OpenStudy (destinyyyy):

Um okay.. This is my first problem with this. And I basically have no clue as to what im doing. And what your saying doesnt follow my examples and you haven't said anything about what ive already solved.

OpenStudy (destinyyyy):

It says to solve for the vertex with x=-b/2a Then plug the answer for x into the problem to get the y-intercept. Factor, then set each factor to 0, that gives the answer x-intercept.

OpenStudy (princeharryyy):

it's a point where you get the maxima or minima

OpenStudy (princeharryyy):